Polly purchased car. She paid ₱ 150,000.00 as down payment and ₱5,500.00 payable at the beginning of each month for 48 months. Money is worth 12% compound monthly.

  • What is the periodic rate?
  • 1%
  • 12%
  • 0.01%
  • 2%
  • How many payments must be made?
  • 12
  • 24
  • 6
  • 48
  • What is the equivalent cash price of the car?
  • ₱210,945.35
  • ₱340,091.59
  • 360,945.35
  • ₱490,091.59
1
Expert's answer
2021-05-10T11:52:12-0400

1.

"periodic \space rate = \frac{stated\space or\space annual\space interest\space rate}{no. \space of \space compounding}"


"=\frac{12\%}{12}=1\%"


2.

48 payments must be made because she needs to make monthly payments at the beginning of each month for 48 months.


3.

"equivalent \space cash price \\= down \space payment + present\space value \space of monthly\space payments"

N = no. of months = 48; I/Y = periodic rate = 1%; PMT = monthly payment = 5,500; FV = future value = 0 > CPT

computing present value we get

PV = 210,945.35

equivalent cash price "= \u20b1150,000 + \u20b1210,945.35 = \u20b1360,945.35"


the equivalent cash price of the car is ₱360,945.35.